Particles A and B carrying charge 3.0nC are at the base corners of an equilateral triangle 2.8 m on a side

Calculate the electric potential energy for the triangular charge distribution?
1
Expert's answer
2021-10-13T09:07:56-0400

The electric potential energy of two point charges is given as follows:


"W = k\dfrac{q_1q_2}{r}"

where "k = 9\times 10^{9}Nm^2/C^2" is the Coulomb constant, "q_1 = q_2 \equiv q = 3\times 10^{-9}C" are the charges (equal in this case), and "r = 2.8m" is the distance between the charges. Thus, obtain:


"W = 9\times 10^{9}\cdot \dfrac{(3\times 10^{-9})^2}{2.8} \approx 3.0\times 10^{-8}J"

ANswer. "3.0\times 10^{-8}J".